10. A square loop (side a) is mounted on a vertical shaft and rotated at angular velocity w. A uniform magnetic field B points to the right. Find the emf for this alternating current generator. 2
Answer #1

Flux F=B*A*Sin theta is angle loop makes with right direction.

Induced emf V= DF/dt=B*A*d(Sinwt)/dt

=B*A*W*Coswt

=B*a2*w*Coswt answer